Todd W. Mummert 1264 Winding Court Mohegan Lake, NY 10547

October 2, 1997

AT&T Wireless Att: Kris D. Ciccarelli

To whom it may concern:

I’d like to obtain cellular service for my wife and I. We’re both IBM employees with the Research Division in Hawthorne, NY.

We’d like the 1-year plan, w/ Nokia 2160 phones ($129/ea). The standard IBM plan ($21 per month, no activation fee, 15 minutes/month, $0.35 peak, $0.30 off-peak) is fine. One optional accessory package ($99) also.

If the above pricing is wrong, please confirm any changes with me at (914) 784-7465 (work) or (914) 245-7443 (home).

With thanks,

Todd W. Mummert